# Dance to Demise – How Massive Stars May Form Dense Circumstellar Shells Before Explosion

### v1

@@ 4.2 Dynamical ejections post shock break-out
The SW prescription of S.-C. Yoon & M. Cantiello (2010) is strictly applicable only to the exponential growth phase of radial pulsations, which subsequently get damped by shocks that develop within the RSG envelope where the radial velocity of the loosely bound layers becomes supersonic. M. Clayton (2018) investigated the problem of dynamical stability of RSG envelopes under such conditions and predicted episodes of repeated mass ejection events which could last for upto $100$-s of years. From their extensive suite of hydrodynamical simulations performed using the MESA code, almost all cases that converged upon a solution after showing such episodes of repeated mass ejections had $log((\mathrm{L}/\mathrm{L}_{\odot})/(\mathrm{M}/\mathrm{M}_{\odot}))\gtrsim 4.1$, for which the following mass loss prescription:
@@
was suggested with a cut-off value of $\log((L/\mathrm{L}_{\odot})/(M/\mathrm{M}_{\odot}))$ between 4.14.154.1-4.15. Hence, we employ this mass loss recipe to model the possibility of such mass ejections occurring post shock-break-out in RSG envelopes when the luminosity to mass ratio of the star exceeds the cut-off value of $\log((L/\mathrm{L}_{\odot})/(M/\mathrm{M}_{\odot}))>4.15$, which is on the conservative side, given the uncertainties in determining the exact value of this cut-off.
@@
Figure 4 shows such a model sequence of initial mass $M_{\mathrm{init}}=16$ M evolved with both the prescriptions given by Equations 2 (for $\alpha=4.35$) and 3 implemented in MESA, with a transition at $\log((L/\mathrm{L}_{\odot})/(M/\mathrm{M}_{\odot}))=4.15$. It can be seen that the onset of the dynamical ejections marks a significant enhancement in the mass loss rates (around an order of magnitude) compared to the superwind phase, which has important implications for the CSM formed by these winds close to the star, as in this model the ejections occur less than $100$ years from CC. Figure 5 shows a set of evolutionary calculations using both the mass loss prescriptions for PDSW (Equation 2) and subsequent dynamical ejections (Equation 3). From our grid of MESA models, we find that only model sequences with initial mass $M_{\rm{init}}\gtrsim 15$ Mevolve to enter the regime for which dynamical ejections are expected (as described in 4.2). Furthermore, depending on the initial mass $M_{\rm{init}}>15$ M, there is a cut-off value of $\alpha$ above which the dynamical ejections lead to elevated mass loss ($\dot{M}\gtrsim 10^{-2}$ M $\rm{yr}^{-1}$).
@@ 4.3 Wind propagation: CSM density, integrated column density
4.3.1 Line-driven acceleration of the wind
Following T. J. Moriya et al. (2018), we consider the line-driven acceleration of the stellar wind which starts with a velocity $v_{o}$ at the launch radius $r=R_{*}$, which is simply the stellar radius. The photons emitted from the luminous RSG surface are absorbed and re-emitted by the wind, which imparts radial momentum leading to gradual acceleration of the wind out to large radii where the wind velocity reaches the terminal speed, $v_{\infty}$. This mechanism of acceleration of the wind is usually approximated the following $\beta$-law for the wind velocity:
@@ 4.3.1 Line-driven acceleration of the wind
where the value of $\beta$ determines the shape of the wind profile. Usually, both $v_{\infty}$ and $\beta$ are obtained through spectral line fitting, with $0.7\lesssim\beta\lesssim 4$ (R.-P. Kudritzki & J. Puls, 2000). In the following comparisons of our results with observed SNe (Section 5), we adopt a constant value of $\beta=1.2$, consistent with hydrodynamical simulations of line-driven winds (M. Curé & I. Araya, 2023).
@@ 4.3.3 Column density along line of sight calculation
with the mean molecular weight $\mu\simeq 1.3$ for hydrogen, and the proton mass $m_{p}=1.6726\times 10^{-24}~gm$. The computed values of the column densities are compared in 5 with available measured values for well-studied SNe (e.g. SNe 2005ip, 2020ywx, 2023ixf), using observed estimates of terminal wind speeds, $v_{\infty}$, required to compute the CSM densities, as per Equation (5).
@@ 5.1 SN 2023ixf
In Figure 6(c), we compare the column density of neutral hydrogen, computed using Equation (7), with measured X-ray values as reported in P. Chandra et al. (2024) and N. A. J. et al. (2025), for which we use the evolution of the shock radius with time as per the estimates of P. Chandra et al. (2024), to convert the epochs into $r$. We observe that, in general, we overpredict the column densities at most epochs where we have observed X-ray measurements. This could arise due to possible global asymmetries in the CSM, as hinted by the rapid spectropolarimetric properties at early times (S. S. Vasylyev et al., 2023), in the ejecta (M. Shrestha et al., 2025) and/or from clumps of material along the line of sight, as indicated by the detection of Fe emission in the X-ray spectra (N. A. J. et al., 2025).
@@
While such effects have been attributed to the ejecta geometry (L. Ferrari et al., 2024) and/or binary interaction (P. D. Michel et al., 2025), regardless of the exact cause of their origins, either scenario would result in underestimation of the integrated column densities along the line of sight in observations, which would make them consistent with our systematically higher model estimates.
5.2 SN 2020ywx
[…]
@@ 5.3 SN 2017hcc
5.4 SN 2005ip
Figure 9(a) shows two sequences of MESA models both with $M_{\text{init}}=16$ M, but with $\alpha=4.36~\rm{and}~4.37$ exhibiting such a sustained period of enhanced mass loss from dynamical ejections following a strong superwind phase with lower mass loss rates between $10^{-4}-10^{-3}$ M $\rm{yr}^{-1}$. The dynamical ejections occur with an average mass loss rate $\gtrsim 10^{-2}$ M $\rm{yr}^{-1}$ over several decades before CC, which leads to $\gtrsim 1$Mof CSM material located close to the progenitor, as seen in Figure 9(b). Our model predictions closely trace the observationally inferred CSM density, computed using a mass loss rate of $1.5\times 10^{-2}$ M $\rm{yr}^{-1}$ (S. Katsuda et al., 2014) and a terminal wind speed of $v_{\infty}=120$ km/s (N. Smith et al., 2009), out to distances of $\sim 5\times 10^{16}$ cm from the progenitor surface which is the location of the forward shock at $\sim 2000$ days post-explosion, assuming the shock velocity $V_{\rm{SH}}\sim t^{-0.1}$ and has a value of $3000~\rm{kms}^{-1}$ at $t=1000$ days (O. D. Fox et al., 2020). At later times, corresponding to larger distances, the mass loss rates are estimated to be much lower at $\sim(2-4)\times 10^{-4}$ M $\rm{yr}^{-1}$ (N. Smith et al., 2009), which is also reflected in our model densities caused due to matter lost in the superwind phase preceding dynamical ejections. The model with $\alpha=4.37$, which has an earlier onset of dynamical ejections at 80~80 yrs from CC, shows this drop in CSM density at a location closer to the observed extent of the dense shell around the progenitor (O. D. Fox et al., 2020).
Figure 9(c) shows the column density profile of neutral hydrogen along the line of sight obtained using the CSM densities according to 4.3.3. The calculated column densities from the model with $\alpha=4.36$ are in close agreement with available estimates from late-time Swift/Chandra X-ray observations of S. Katsuda et al. (2014) and O. D. Fox et al. (2020), which are converted into radial coordinates from post-explosion times using the time evolution of the shock wavefront, as mentioned above.
@@ 5.5 SN 1998S
SN 1998S is considered to be the prototypical Type-IIn SN, with a mass loss rate estimate of $6\times 10^{-3}$ M $\rm{yr}^{-1}$ during the last $\sim 15$ yrs before CC and a much lower rate at earlier times (I. Shivvers et al., 2015), though J. Mauerhan & N. Smith (2012) showed substantial matter present as far out as $r\sim 2\times 10^{17}$ cm from late-time spectroscopic observations $14$ yrs post-explosion. Previously, the terminal velocity of the wind and the mass loss rate of the progenitor had been estimated to be 405040-50 km/s and 121-2 M $\rm{yr}^{-1}$ respectively (A. Fassia et al., 2001; D. Pooley et al., 2002). Based on these studies, we use the following observed mass loss rate profile:
@@
along with a terminal wind speed $v_{\infty}=40$ $v_{0}=0.1 km/s$ and $R_{*}=1.1\times 10^{14}~cm$ to calculate the observationally inferred CSM density as per Equation 5.
@@
In left panel of Figure 10, we present our $18$ $\alpha=2.83 M$ESA model sequences which shows enhanced mass loss rates ($\dot{M}>10^{-4}$ M $\rm{yr}^{-1}$) due to superwinds from $\sim 10^{3.5}$ yrs up to $\sim 16.22$ yrs before CC, after which there is a sudden increase in mass loss rate ($\sim 10^{-2}$ M $\rm{yr}^{-1}$) from episodes of dynamical ejections (as described in Section 4.2), until $\sim 1.48$ yr prior to CC. These properties are similar to the observed estimates for the mass loss of the progenitor of SN 1998S as discussed above, with mass loss rates as given by Equation 8 within a factor of $\sim 2$ of those in these MESA model sequences.
The right hand panel of Figure 10 compares the CSM density predicted from the MESA models using Equation 5 with the observationally inferred values. The ”inner” CSM ($r<10^{15}$ cm) predictions differ from the observational estimated densities by up to an order of magnitude, which may arise due to the constant values assumed for the observed mass loss rate, stellar radius and the launch velocity of the wind at the surface of the star. Other evolutionary scenarios invoked to explain the observed light curve of SN 1998S, that requires a dense CSM nearby, include merger of a massive RSG with a compact object (S. L. Schrøder et al., 2020). However, the presence of two distinct shells as described above poses tension with this merger-driven explosion scenario, though episodic mass ejections may still occur during the common-envelope (CE) phase in such systems (M. Clayton et al., 2017).
@@ 6.1 Observational implications for the RSG problem
Our MESA evolutionary sequences with the PDSW and subsequent dynamical ejections for initial masses $\lesssim 18$ M, presented in 4, undergo late-stage heavy mass loss, creating the dense CSM around the progenitor star as is required to explain observations of the Type II SNe shown in 5. S.-C. Yoon & M. Cantiello (2010) suggested that, considering their PDSW mass loss prescription, their exists a critical mass, $M_{\rm{crit}}\sim 19-20$ M above which RSGs lose a large fraction of their H-rich envelopes prior to CC, without producing Type IIP SNe, reconciling with the so-called missing RSG problem inferred from observations of RSGs (S. J. Smartt et al., 2009; S. J. Smartt, 2015; B. Davies & E. R. Beasor, 2020; C. S. Kochanek, 2020; E. R. Beasor et al., 2025; Q. Fang et al., 2025). We performed some calculations with higher initial mass in the above range, most of which led to runaway mass loss due to shock-driven mass ejections shortly following the onset of PDSW, in the very early stages of core-carbon burning. We could not evolve these runs further due to numerical convergence issues, but such runaway events also favor the scenario in which most of the H-rich envelope lost much before CC, in agreement with the observed lack of higher mass RSGs associated with Type IIP SNe.
@@ 7 Conclusions
This work presents a comprehensive, end-to-end investigation of how pulsating RSGs can create a dense, possibly dusty, CSM through enhanced mass loss episodes in the final years before CC, which shapes the multi-wavelength appearance of the ensuing SN explosion. From a dense grid of MESA models of varying $M_{\rm{init}}=12-20$ M  and PDSW strength (in terms of the parameter $\alpha=2-5$, as in S.-C. Yoon & M. Cantiello 2010), as well as subsequent shock-driven dynamical ejections (M. Clayton, 2018), the mass loss histories of the model sequences are coupled to an accelerated $\beta$-law wind profile to construct spherically symmetric CSM density and column-density profiles that are directly compared to multi-epoch observations of SNe 1998S, 2005ip, 2017hcc, 2020ywx, and 2023ixf. We find decent agreement (within a factor of few) between the observations and our model predictions for $M_{\rm{init}}>15$ M  and $\alpha>2$ which give enhanced mass loss rates ($\dot{M}\sim 10^{-4}-10^{-2}$ M yr-1) in the final decades to centuries before explosion. Such late phase heavy mass loss produces sufficiently dense CSM close to the SN progenitor to explain early-time observations (e.g. using flash-spectroscopy, X-rays and radio wavelengths) as well as late epoch estimates of X-ray column densities (Section 5). Our results quantitatively demonstrate that single star evolutionary pathways with pulsation-driven mass loss and dynamical ejections can account for the observed diversity in the properties of Type II SN environments.
In future, missing physics such as effects of varying rotation and metallicity as well as binary interaction needs to be incorporated to improve the realism of our models, all of which are expected to have broader implications both for massive star evolution as well as their final fates in terms of the nature of the eventual SN. Also, the pulsation properties of other classes of massive stars such as WR stars (L. Grassitelli et al., 2016, 2018) and LBVs (J. A. Guzik et al., 1999; S. Justham et al., 2014) deserve investigation as they are supposed to be prime candidates for progenitors of both Type Ib/c and superluminous SNe. In addition, further improvements in our understanding of the interplay of radial pulsation and convective motions in RSG envelope through multidimensional simulations and its role in driving mass loss will have important implication for SN modeling at large and their impact on cosmic evolution and dust formation ranging from our local environment to the high redshift universe in the era of large-scale surveys like the JWST and the LSST.

### v2

@@ 4.2 Dynamical ejections post shock break-out
The SW prescription of S.-C. Yoon & M. Cantiello (2010) is strictly applicable only to the exponential growth phase of radial pulsations, which subsequently get damped by shocks that develop within the RSG envelope where the radial velocity of the loosely bound layers becomes supersonic. M. Clayton (2018) investigated the problem of dynamical stability of RSG envelopes in the shock-dominated regime, in terms of the pressure-weighted, volume averaged first adiabatic exponent (see Section 3.4 in  M. Clayton et al., 2017) whose value less than $4/3$ at any point within the stellar envelope is indicative of instability to dynamical ejections from that point outwards (A. Lobel, 2001) - commonly occurring in regions above the hydrogen and helium ionization zones. Detailed shock-resolving hydrodynamic MESA models of M. Clayton (2018) showed launch of ejections upon the surface breakout of compression shocks followed by sufficiently fast expansion that raises material on to an escape trajectory. M. Clayton (2018) implemented a wind-mass loss scheme in order to excise any such contiguous unbound sub-surface layer of mass $m$ (with radial velocity exceeding the local escape velocity), by applying a mass loss rate of $100m$ $\rm{yr}^{-1}$ corresponding to a time-scale which is at least one order of magnitude below both the dynamical and thermal time-scales of the star. A subset of their extensive suite of hydrodynamic simulations of such shock-driven mass loss resulted in episodes of repeated mass ejection events lasting for upto $100$-s of years. Almost all cases that converged upon a solution after showing such episodes of repeated mass ejections had $\log((\mathrm{L}/\mathrm{L}_{\odot})/(\mathrm{M}/\mathrm{M}_{\odot}))\gtrsim 4.1$, for which M. Clayton (2018) extracted time-averaged mass loss rates to propose the following mass loss prescription:
@@
above a cut-off value of $\log((L/\mathrm{L}_{\odot})/(M/\mathrm{M}_{\odot}))$ between 4.14.154.1-4.15. Hence, we employ this mass loss recipe to model the possibility of such repeating mass ejections occurring post shock break-out in pulsationally unstableRSG envelopes when the luminosity to mass ratio of the star exceeds the cut-off value of $\log((L/\mathrm{L}_{\odot})/(M/\mathrm{M}_{\odot}))=4.15$, which is on the conservative side, given the uncertainties in determining the exact value of this cut-off (M. Clayton, 2018). We do not consider any other mass loss mechanism in this regime, so as to avoid duplicating effects of any similar mechanism(s) triggering mass loss in the final years leading to CC. Furthermore, the above recipe does not account for other dynamically unstable scenarios which lead to single ejection episodes, also identified in the simulations of M. Clayton (2018), that occur on a much shorter timescale to be accounted for in our evolutionary calculations.
@@
Figure 4 shows such a model sequence of initial mass $M_{\mathrm{init}}=16$ M evolved with both the prescriptions given by Equations 2 (for $\alpha=4.35$) and 3 implemented in MESA, with a transition at $\log((L/\mathrm{L}_{\odot})/(M/\mathrm{M}_{\odot}))=4.15$. It can be seen that the onset of the dynamical ejections marks a significant enhancement in the mass loss rates (around an order of magnitude) compared to the superwind phase, which has important implications for the CSM formed by these winds close to the star, as in this model the ejections occur less than $100$ years from CC. Figure 5 shows a set of evolutionary calculations using both the mass loss prescriptions for PDSW (Equation 2) and subsequent dynamical ejections (Equation 3). From our grid of MESA models, we find that only model sequences with initial mass $M_{\rm{init}}>15$ M  evolve to enter the regime for which dynamical ejections are expected (as described in 4.2) for $\alpha=2-5$. Furthermore, depending on the initial mass, there is a cut-off value of $\alpha$ above which the dynamical ejections lead to elevated mass loss ($\dot{M}\gtrsim 10^{-2}$ M $\rm{yr}^{-1}$).
@@ 4.3 Wind propagation: CSM density, integrated column density
4.3.1 Acceleration of the RSG wind
Following T. J. Moriya et al. (2018), we consider acceleration of the stellar wind – which starts with a velocity $v_{o}$ at the launch radius $r=R_{*}$, which is is assumed to be just the stellar radius – out to large radii where the wind velocity gradually reaches a constant terminal speed, $v_{\infty}$. Since the exact mechanism of acceleration of the wind remains unknown, the kinematics is usually parametrized in form of the following simple $\beta$-law for the wind velocity (H. J. G. L. M. Lamers & J. P. Cassinelli, 1999):
@@ 4.3.1 Acceleration of the RSG wind
where the value of $\beta$ determines the shape of the wind profile. Usually, both $v_{\infty}$ and $\beta$ are obtained through spectral line fitting for which $0.7\lesssim\beta\lesssim 4$ has been used in the literature (R.-P. Kudritzki & J. Puls, 2000). RSGs known to experience slower wind acceleration as compared to hotter (OB) stars i.e. $\beta\gtrsim 1$ (R. Baade et al., 1996; P. D. Bennett, 2010). For the following comparisons of our results with CSM properties of observed SNe (Section 5), we adopt a constant value of $\beta=1.2$, consistent with hydrodynamical simulations (M. Curé & I. Araya, 2023) as well as observations of single RSGs (P. D. Bennett 2010; see Appendix A) while the values of the terminal velocity, $v_{\infty}$, are used specific to each SN as inferred from fitting of spectral lines.
@@ 4.3.3 Column density along line of sight calculation
with the mean molecular weight $\mu\simeq 1.3$ for hydrogen, and the proton mass $m_{p}=1.6726\times 10^{-24}~gm$. The computed values of the column densities are compared in 5 with available measured values for well-studied SNe (1998S, 2005ip, 2020ywx, 2023ixf), using observed estimates of terminal wind speeds, $v_{\infty}$, required to compute the CSM densities, as per Equation (5).
@@ 5.1 SN 2023ixf
In Figure 6(c), we compare the column density of neutral hydrogen, computed using Equation (7), with measured X-ray values as reported in P. Chandra et al. (2024) and N. A. J. et al. (2025), for which we use the evolution of the shock radius with time as per the estimates of P. Chandra et al. (2024), to convert the epochs into $r$. We observe that, in general, we overpredict the column densities at most epochs where we have observed X-ray measurements. This could arise due to possible global asymmetries in the CSM, as hinted by the rapid spectropolarimetric properties at early times in the ejecta (S. S. Vasylyev et al., 2023; M. Shrestha et al., 2025) and/or from clumps of material along the line of sight, as indicated by the detection of Fe emission in the X-ray spectra (N. A. J. et al., 2025). While such effects have been attributed to the ejecta geometry (L. Ferrari et al., 2024) and/or binary interaction (P. D. Michel et al., 2025), regardless of the exact cause of their origins, either scenario would result in underestimation of the integrated column densities along the line of sight in observations, which would make them consistent with our systematically higher model estimates.
5.2 SN 2020ywx
[…]
@@ 5.2 SN 2020ywx
effects or line-of-sight effects influencing the column density measurements of R. Baer-Way et al. (2025), as discussed for SN 2023ixf (P. Chandra et al., 2024; N. A. J. et al., 2025).
@@ 5.3 SN 2017hcc
5.4 SN 2005ip
Figure 9(a) shows two sequences of MESA models both with $M_{\text{init}}=16$ M, but with $\alpha=4.36~\rm{and}~4.37$ exhibiting such a sustained period of enhanced mass loss from dynamical ejections following a strong superwind phase with lower mass loss rates between $10^{-4}-10^{-3}$ M $\rm{yr}^{-1}$. The dynamical ejections occur with an average mass loss rate $\gtrsim 10^{-2}$ M $\rm{yr}^{-1}$ over several decades before CC, which leads to $\gtrsim 1$Mof CSM material located close to the progenitor, as seen in Figure 9(b). Our model predictions closely trace the observationally inferred CSM density, computed using a mass loss rate of $1.5\times 10^{-2}$ M $\rm{yr}^{-1}$ (S. Katsuda et al., 2014) and a terminal wind speed of $v_{\infty}=120$ $\rm{km~s}^{-1}$ (N. Smith et al., 2009), out to distances of $\sim 5\times 10^{16}$ cm from the progenitor surface which is the location of the forward shock at $\sim 2000$ days post-explosion, assuming the shock velocity $V_{\rm{SH}}\sim t^{-0.1}$ and has a value of $3000~\rm{kms}^{-1}$ at $t=1000$ days (O. D. Fox et al., 2020). At later times, corresponding to larger distances, the mass loss rates are estimated to be much lower at $\sim(2-4)\times 10^{-4}$ M $\rm{yr}^{-1}$ which is also reflected in our model densities caused due to matter lost in the superwind phase preceding dynamical ejections. The model with $\alpha=4.37$, which has an earlier onset of dynamical ejections at 80~80 yrs from CC, (N. Smith et al., 2009), shows this drop in CSM density at a location closer to the observed extent of the dense shell around the progenitor (O. D. Fox et al., 2020).
@@ 5.4 SN 2005ip
Figure 9(c) (N. Smith et al., 2009), shows the column density profile of neutral hydrogen along the line of sight obtained using the CSM densities according to 4.3.3. The calculated column densities from the model with $\alpha=4.36$ are in close agreement with available estimates from late-time Swift/Chandra X-ray observations of S. Katsuda et al. (2014) and O. D. Fox et al. (2020), which are converted into radial coordinates from post-explosion times using the time evolution of the shock wavefront, as mentioned above.
@@ 5.5 SN 1998S
SN 1998S is considered to be the prototypical Type-IIn SN, with a mass loss rate estimate of $6\times 10^{-3}$ M $\rm{yr}^{-1}$ during the last $\sim 15$ yrs before CC and a much lower rate at earlier times (I. Shivvers et al., 2015), though J. Mauerhan & N. Smith (2012) showed substantial matter present as far out as $r\sim 2\times 10^{17}$ cm from late-time spectroscopic observations $14$ yrs post-explosion. Previously, the terminal velocity of the wind and the mass loss rate of the progenitor had been estimated to be 405040-50 km/s and $(1-2)\times 10^{-4}$ M $\rm{yr}^{-1}$ respectively (A. Fassia et al., 2001; D. Pooley et al., 2002). Based on these studies, we use the following observed mass loss rate profile:
@@
along with a terminal wind speed $v_{\infty}=40$ $\rm{km~s}^{-1}$$v_{0}=0.1$ km/s and $R_{*}=1.1\times 10^{14}~cm$ to calculate the observationally inferred CSM density as per Equation 5.
@@
In top panel of Figure 10, we present our $18$ $\alpha=2.83 M$ESA model sequences which shows enhanced mass loss rates ($\dot{M}>10^{-4}$ M $\rm{yr}^{-1}$) due to superwinds from $\sim 10^{3.5}$ yrs up to $\sim 16.22$ yrs before CC, after which there is a sudden increase in mass loss rate ($\sim 10^{-2}$ M $\rm{yr}^{-1}$) from episodes of dynamical ejections (as described in Section 4.2), until $\sim 1.48$ yr prior to CC. These properties are similar to the observed estimates for the mass loss of the progenitor of SN 1998S as discussed above, with mass loss rates as given by Equation 8 within a factor of $\sim 2$ of those in these MESA model sequences.
The bottom panel of Figure 10 compares the CSM density predicted from the MESA models using Equation 5 with the values inferred from the observed mass loss rate profile given by Equation (8). The “inner” CSM ($r<10^{15}$ cm) predictions closely resemble observationally estimated density profile for $r\lesssim 7\times 10^{14}~\mathrm{cm}$. Other evolutionary scenarios invoked to explain the observed light curve of SN 1998S, that requires a dense CSM nearby, include merger of a massive RSG with a compact object (S. L. Schrøder et al., 2020). However, the presence of two distinct shells as described above poses tension with this merger-driven explosion scenario, though episodic mass ejections may still occur during the common-envelope (CE) phase in such systems (M. Clayton et al., 2017).
@@ 6.1 Observational implications for the RSG problem
Our MESA evolutionary sequences with the PDSW and subsequent dynamical ejections for initial masses $\lesssim 18$ M, presented in Section 4, undergo late-stage heavy mass loss, creating the dense CSM around the progenitor star as is required to explain observations of the Type II SNe shown in Section 5. S.-C. Yoon & M. Cantiello (2010) suggested that, considering their PDSW mass loss prescription, their exists a critical mass, $M_{\rm{crit}}\sim 19-20$ M above which RSGs lose a large fraction of their H-rich envelopes prior to CC, without producing Type IIP SNe, reconciling with the so-called missing RSG problem inferred from observations of RSGs (S. J. Smartt et al., 2009; S. J. Smartt, 2015; B. Davies & E. R. Beasor, 2020; C. S. Kochanek, 2020; E. R. Beasor et al., 2025; Q. Fang et al., 2025). We performed some calculations with higher initial mass in the above range, most of which led to runaway mass loss due to shock-driven mass ejections shortly following the onset of PDSW, in the very early stages of core-carbon burning. We could not evolve these runs further due to numerical convergence issues, but such runaway events also favor the scenario in which most of the H-rich envelope lost much before CC, in agreement with the observed lack of higher mass RSGs associated with Type IIP SNe.
@@ 7 Conclusions
This work presents a comprehensive, end-to-end investigation of how pulsating RSGs can create a dense, possibly dusty, CSM through enhanced mass loss episodes in the final years before CC, which shapes the multi-wavelength appearance of the ensuing SN explosion. From a dense grid of MESA models of varying $M_{\rm{init}}=12-20$ M  and PDSW strength (in terms of the superwind parameter $\alpha=2-5$, as in S.-C. Yoon & M. Cantiello 2010), as well as subsequent shock-driven dynamical ejections (M. Clayton, 2018), the mass loss histories of the model sequences are coupled to an accelerated $\beta$-law wind profile to construct spherically symmetric CSM density and column-density profiles that are directly compared to multi-epoch observations of SNe 1998S, 2005ip, 2017hcc, 2020ywx, and 2023ixf. We find decent agreement (within a factor of few) between the observationally inferred values and our model results for $M_{\rm{init}}>15$ M  and $\alpha>2$ which give enhanced mass loss rates ($\dot{M}\sim 10^{-4}-10^{-2}$ M yr-1) in the final decades to centuries before explosion. Such late phase heavy mass loss produces sufficiently dense CSM close to the SN progenitor to explain early-time observations (e.g. using flash-spectroscopy, X-rays and radio wavelengths) as well as late epoch estimates of X-ray column densities (Section 5). Our results quantitatively demonstrate that single star evolutionary pathways with pulsation-driven mass loss and consequent dynamical ejections can account for the observed diversity in the properties of Type II SN environments.
In future, missing physics such as effects of varying rotation and metallicity as well as binary interaction needs to be incorporated to improve the realism of our models, all of which are expected to have broader implications both for massive star evolution as well as their final fates in terms of the nature of the eventual SN. Also, the pulsation properties of other classes of massive stars such as WR stars (L. Grassitelli et al., 2016, 2018) and LBVs (J. A. Guzik et al., 1999; S. Justham et al., 2014) deserve investigation as they are supposed to be prime candidates for progenitors of both Type Ib/c and superluminous SNe. In addition, further improvements in our understanding of the interplay of radial pulsation and convective motions in RSG envelope through multidimensional simulations (J. A. Goldberg et al., 2022, 2025) and its role in driving mass loss are critical towards advances in the predictive power of evolutionary models by incorporation self-consistent treatment of stellar winds through physical mechanisms such as pulsations and shocks, as compared to the empirical approach adopted in this study and extant works in the literature (S.-C. Yoon & M. Cantiello, 2010; Moriya, Takashi J. & Langer, Norbert, 2015; V. A. Bronner et al., 2025). This will have important implications for SN modeling at large and their impact on cosmic evolution and dust formation ranging from our local environment to the high redshift universe in the era of large-scale surveys like the JWST and the LSST.